1Problem 1 . . Long-term care (LTC) policies are separated into two groups: professional and home care. Dave examines professi
Answer #1

For positive integer n, the probability that Atharv examines exactly n policies is (1-0.3)n-1 *0.3 =0.7n-1 *0.3

the probability that Dave examines more than n policies is (1-0.25)n =0.75n

since they are mutually independent

The probability that Artharv examines more policies than Dave is

Σ0.7-! * 0.3 * 0.75 7t  ,n=1,2,....

= 0.3/0.7 〉 0.7 * 0.75 n=1

= 0.43 Σ 0.525 n=1 7l

=0.43* 0.525/(1-0.525)

=0.48